Solve the following quadratic equation by using formula
6x²-x-12=0
Can I be shown this in detail please

Answer by stanbon(75887)

Solve the following quadratic equation by using formula
6x²-x-12=0
-----
Formula:: x = [-b +- sqrt(b^2-4ac)]/(2a)
a = 6 ; b = -1 ; c = -12
----------------------------------
x = [1 +- sqrt(1-4*6*-12)]/(12)
-----
x = [1 +- sqrt(1+sqrt(288)/12
-----
x = [1 +- 17]/12
----
x = 18/12 or x = -16/12
----
x = 3/2 or x = -4/3
------------------------------
Cheers,
Stan H.
------------------